# Commute Graph Neural Networks

### v1

@@ 4.2 Similarity-based Graph Rewiring
Both the fundamental matrix defined in Eq. 3 and Weighted DiLap requires 3.1 to ensure the existence and uniqueness of the Perron vector $\pi$, conditions that are not universally met in general graphs. To fulfill the irreducibility and aperiodicity assumptions,  Tong et al. (2020a) introduce a teleporting probability uniformly distributed across all nodes. This method, inspired by PageRank (Page et al., 1999), amends the transition matrix to $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, where $\gamma\in(0,1)$. $\mathbf{P}_{pr}$ allows for the possibility that a random walker may choose a non-neighbor node for the next step with a probability of $\frac{1-\gamma}{N}$. This adjustment ensures that $\mathbf{P}_{pr}$ is irreducible and aperiodic, so it has a unique $\pi$. However, this approach leads to a complete graph represented by a dense matrix $\mathbf{P}_{pr}$, posing significant challenges for subsequent computational processes.
@@
This rewiring approach satisfies 3.1 and maintains graph sparsity. Additionally, adding edges between nodes with similar features only minimally alters the overall semantics of the original graph. Based on $\widetilde{G}$ and its corresponding $\widetilde{\mathbf{P}}$ and $\widetilde{\mathbf{\Pi}}$, we have the deterministic Weighted DiLap $\widetilde{\mathcal{T}}$.
@@ 4.3 From DiLap to Commute Time
The proof is given in Section A.2. Leveraging Lemma 4.1, we can further compute the hitting times and commute times in terms of $\widetilde{\mathcal{T}}$ with the following theorem.
@@
For a detailed proof of this theorem, refer to Theorem 3 in (Li & Zhang, 2012), where the corresponding term should be replaced with $\mathcal{R}$. Then we can derive the matrix forms of the hitting time $\mathcal{H}$ and commute time $\mathcal{C}$ as per Eq. 9:
@@
where $\odot$ denotes Hadamard product, and $\otimes$ is outer product. $\mathcal{H}_{ij}$ and $\mathcal{C}_{ij}$ correspond to the hitting and commute time from $v_{i}$ to $v_{j}$ respectively. The computation of commute times via DiLap, in contrast to the method delineated in Theorem 3.2, is primarily motivated by efficiency concerns. Specifically, Eq. 3 necessitates the inversion of a dense matrix with complexity $\mathcal{O}(N^{3})$, whereas our DiLap-based method hinges on computing the pseudoinverse of a sparse matrix $\mathcal{R}$. The pseudoinverse of $\mathcal{R}$ can be efficiently determined using SVD. Given the sparse nature of $\mathcal{R}$, we can employ well-established techniques such as the randomized truncated SVD algorithm (Halko et al., 2011; Cai et al., 2023), which takes advantage of sparsity, to reduce the time complexity to $\mathcal{O}(q|E|)$, where $|E|$ denotes the number of edges reflecting the sparsity (See Section A.3). Next, we present Commute Graph Neural Networks (CGNN) based on $\mathcal{C}$.
@@ 4.4 CGNN
where $\mathrm{Agg}_{\text{in}}^{(\ell)}(\cdot)$ and $\mathrm{Agg}_{\text{out}}^{(\ell)}(\cdot)$ are mean aggregation functions with different feature transformation weights, and $\mathrm{Comb}^{(\ell)}(\cdot)$ is a mean operator. Within each layer, the influence of $v_{j}$ on the central node $v_{i}$ is modulated by the commute-time-based proximity $\widetilde{\mathcal{C}}$ based on the edge directionality. We present the pseudocode and complexity of CGNN in Algorithm 1.
5 Experiments
We conduct node classification experiments on five digraph datasets. Experimental details and data statistics are provided in Section C.1 and Section C.2. We provide a performance comparison with 12 baselines including 1) General GNNs: GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017); 2) Non-local GNNs: APPNP (Klicpera et al., 2019), MixHop (Abu-El-Haija et al., 2019), GPRGNN (Chien et al., 2021), and GCNII (Ming Chen et al., 2020); 3) Digraph NNs: DGCN (Tong et al., 2020b), DiGCN (Tong et al., 2020a), MagNet (Zhang et al., 2021), DiGCL (Tong et al., 2021), and DirGNN (Rossi et al., 2023). For all baselines, we apply both the symmetrized and asymmetric adjacency matrix for experiments. The results reported are the better of the two results.
Table 1: Node classification results. Accuracy (%) with standard deviation for 10 runs. We highlight/underline the best/second best method. For general GNN and non-local GNN baselines, we conduct experiments on both symmetrized versions and their directed counterparts, reporting better results from these two settings.
Method Squirrel Chameleon Citeseer CoraML AM-Photo GCN 52.43$\pm$2.01 67.96$\pm$1.82 66.03$\pm$1.88 70.92$\pm$0.39 88.52$\pm$0.47 GAT 40.72$\pm$1.55 60.69$\pm$1.95 65.58$\pm$1.39 72.22$\pm$0.57 88.36$\pm$1.25 GraphSAGE 41.61$\pm$0.74 62.01$\pm$1.06 66.81$\pm$1.38 74.16$\pm$1.55 89.71$\pm$0.57 APPNP 51.91$\pm$0.56 45.37$\pm$1.62 66.90$\pm$1.82 70.31$\pm$0.67 87.43$\pm$0.98 MixHop 43.80$\pm$1.48 60.50$\pm$2.53 56.09$\pm$2.08 65.89$\pm$1.50 87.17$\pm$1.34 GPRGNN 50.56$\pm$1.51 66.31$\pm$2.05 61.74$\pm$1.87 73.31$\pm$1.37 90.23$\pm$0.34 GCNII 38.47$\pm$1.58 63.86$\pm$3.04 58.32$\pm$1.93 64.84$\pm$0.71 83.40$\pm$0.79 DGCN 37.16$\pm$1.72 50.77$\pm$3.31 66.37$\pm$1.93 75.02$\pm$0.50 87.74$\pm$1.02 DiGCN 33.44$\pm$2.07 50.37$\pm$4.31 64.99$\pm$1.72 77.03$\pm$0.70 88.66$\pm$0.51 MagNet 39.01$\pm$1.93 58.22$\pm$2.87 65.04$\pm$0.47 76.32$\pm$0.10 86.80$\pm$0.65 DiGCL 35.82$\pm$1.73 56.45$\pm$2.77 67.42$\pm$0.14 77.53$\pm$0.14 89.41$\pm$0.11 DirGNN 75.19$\pm$1.26 79.11$\pm$2.28 66.57$\pm$0.74 75.33$\pm$0.32 88.09$\pm$0.46 CGNN 77.20$\pm$2.42 79.54$\pm$1.82 70.27$\pm$0.66 77.02$\pm$0.67 90.41$\pm$0.32
5.1 Overall Results
Table 1 reports the node classification results across five digraph datasets. Our method CGNN achieves new state-of-the-art results on 4 out of 5 datasets, and comparable results on the remaining one, validating the superiority of CGNN. We provide more observations as follows. Firstly, while non-local GNNs have the potential to cover the commute paths between adjacent nodes by stacking multiple layers, they do not consistently outperform general, shallow GNN models. It suggests that coarsely aggregating all nodes in commute paths is ineffective. The reason is that deeper models may aggregate excessive irrelevant information for the central node. Our goal is to encode mutual relationships between adjacent nodes by considering their commute times. Aggregating all nodes along the entire path introduces excessive information about other nodes unrelated to the direct relationship between the target nodes. Secondly, GNNs tailored for digraphs do not seem to bring substantial gains. Our results show that with careful hyper-parameter tuning, general GNNs can achieve results comparable to, or even better than, those tailored for digraphs, as evidenced in the Squirrel, Chameleon, and AM-Photo datasets. Thirdly, CGNN achieves state-of-the-art results on both homophilic and heterophilic digraph benchmarks. Notably, DirGNN also performs comparably on heterophilic graphs (Squirrel and Chameleon), supporting the findings of Rossi et al. (2023) that distinguishing edge directionality during message passing enables the central node to adaptively balance information flows from both heterophilic and homophilic neighbors, effectively mitigating the impact of heterophily. Moreover, CGNN, an enhanced version of DirGNN, further improves performance on these graphs by effectively incorporating commute times to refine the strength of relationships between nodes, enhancing model robustness under heterophily.
(a) Squirrel
(b) AM-Photo
Figure 3: Accuracy vs. running time on Squirrel and AM-Photo.
5.2 Efficiency Comparsion
5.3 Component Analysis
Comparison between graph rewiring and PPR.
In Section 4.2, we construct a rewired graph $\widetilde{G}$ based on feature similarity to guarantee the irreducibility and aperiodicity. On the other hand, In contrast, the classic PageRank transition matrix, defined as $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, achieves a similar objective but results in a completely connected graph $G_{pr}$. However, this approach tends to overlook the sparse structure of the original graph, which may alter the semantic information in the graph. Additionally, computing commute times using a dense transition matrix incurs a high computational cost. To validate the effectiveness of the rewiring approach over the PPR method, we conduct an experiment where $\widetilde{G}$ is replaced with $G_{pr}$ in the computation of commute-time-based proximity. We denote this variant as ‘CGNNppr’ and the results of accuracy and efficiency are reported in Table 3. The findings reveal that the PPR approach is suboptimal in terms of both accuracy and efficiency, thereby underscoring the effectiveness of our rewiring-based approach.
Table 3: Accuracy and running time (s) of CGNN and CGNNppr.
Squirrel Chameleon Citeseer CoraML AM-Photo Method Acc. Time Acc. Time Acc. Time Acc. Time Acc. Time CGNN 77.20 102.93 79.54 112.24 80.27 92.26 77.02 127.15 90.41 121.80 CGNNppr 68.37 257.84 71.69 253.05 78.59 137.82 76.23 192.09 88.52 203.04
Figure 4: Impact of directed structure.
Directed vs. Undirected.
To validate the critical role of directed structures in our model, we transform all directed edges into undirected ones by adding their reverse counterparts. This process results in a symmetric adjacency matrix, denoted as $\mathbf{A}_{\mathrm{sym}}$. Subsequently, the commute time is calculated based on the transition matrix derived from $\mathbf{A}_{\mathrm{sym}}$. We refer this variant as ‘CGNNsym’. Fig. 4 shows the accuracy of CGNN and CGNNsym on three datasets. We find that edge direction can significantly influence the prediction accuracy for our model.
6 Conclusion
In this work, we introduce the Commute Graph Neural Network (CGNN) to integrate commute time into GNNs for digraphs. To this end, we propose DiLap, a new Laplacian formulation for digraphs, along with a rapid computational method for determining commute times. By integrating commute times into GNN message passing through neighbor weighting, CGNN effectively leverages the path asymmetry in digraphs, thus improving node representation learning. Our comprehensive experiments demonstrate the superior performance of CGNN over existing methods.
Appendix A Proofs and Derivations
A.1 Derivation of DiLap T
Since the digraph is unweighted, then $\mathbf{P}_{ij}=\frac{1}{d^{\text{out}}_{i}}$ and $\frac{1}{d^{\text{out}}_{i}}\mathbf{A}_{i,:}=\mathbf{P}_{i,:}$. We have:
$(\mathbf{T}s)_{i}$ $=\sum_{v_{j}\in\mathcal{N}^{\text{out}}_{i}}\mathbf{P}_{ij}(s_{i}-s_{j})$ (12)
$=\frac{1}{d^{\text{out}}_{i}}s_{i}-\frac{1}{d^{\text{out}}_{i}}\sum_{v_{j}\in\mathcal{N}^{\text{out}}_{i}}s_{j}$
$=(\mathbf{D}^{-1}s)_{i}-\frac{1}{d^{\text{out}}_{i}}\mathbf{A}_{i,:}s$
$=(\mathbf{D}^{-1}s)_{i}-(\mathbf{P}s)_{i}$
$=((\mathbf{D}^{-1}-\mathbf{P})s)_{i}.$
Thus, DiLap $\mathbf{T}$ can be represented as $\mathbf{D}^{-1}-\mathbf{P}$.
A.2 Proof of Lemma 4.1
Proof.
Since the transition matrix $\widetilde{\mathbf{P}}$ is row-stochastic, it follows that $\widetilde{\mathbf{P}}^{t}\mathbf{J}=\mathbf{J}$. In light of Eq. 2 and considering that $\pi$ is stochastic, we have $\mathbf{Z}\mathbf{J}=\bm{0}_{n\times n}$. Let $\mathcal{K}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\widetilde{\mathcal{T}}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}(\widetilde{\mathbf{D}}^{-1}-\widetilde{\mathbf{P}})\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}$, $\mathcal{J}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}\mathbf{J}\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}$, and $\mathcal{Z}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}\mathbf{Z}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}$, incorporating these into Eq. 3, we have:
$\mathcal{Z}+\mathcal{J}=(\mathcal{K}+\mathcal{J}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})^{-1}.$ (13)
By post-multiplying Eq. 13 from the right by $(\mathcal{K}+\mathcal{J}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})$, we have:
$\mathbf{I}$ $=(\mathcal{Z}+\mathcal{J})(\mathcal{K}+\mathcal{J}-\tilde{\mathbf{D}}_{c}^{-1}+\mathbf{I})$ (14)
$=\mathcal{Z}\mathcal{K}+\mathcal{Z}\mathcal{J}-\mathcal{Z}\widetilde{\mathbf{D}}^{-1}+\mathcal{Z}+\mathcal{J}\mathcal{K}+\mathcal{J}^{2}-\mathcal{J}\widetilde{\mathbf{D}}^{-1}+\mathcal{J}.$
Since $\mathcal{J}^{2}=\mathcal{J}$, $\mathcal{J}\mathcal{K}=\mathcal{J}(\widetilde{\mathbf{D}}^{-1}-\mathbf{I})$, and $\mathcal{Z}\mathcal{J}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}\mathbf{Z}\mathbf{J}\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}=\bm{0}_{n\times n}$, Eq. 14 can be simplified to:
$\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})=\mathbf{I}-\mathcal{J}.$ (15)
Similarly, by multiplying from the left, we establish that $(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}=\mathbf{I}-\mathcal{J}$. As $\pi^{\top}\mathbf{Z}=\bm{0}$, we have $\mathcal{J}\mathcal{Z}=0$. Thus, $\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}=\mathcal{Z}$. Furthermore, it can be straightforwardly inferred that $(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{J}=0$, then have $(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})=(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})$. Besides, considering the symmetry of the left part of Eq. 15, we have $(\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I}))^{\top}=\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})$. Similarly, $((\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z})^{\top}=(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}$. These derivations satisfy the sufficient conditions for the Moore–Penrose pseudoinverse, such that
$\mathcal{Z}=(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})^{\dagger}.$ (16)
Finally, recovering $\mathcal{Z}$ and $\mathcal{K}$, which concludes the proof. ∎
A.3 SVD for $\mathcal{R}^{\dagger}$
Given a matrix $\mathcal{R}\in\mathbb{R}^{N\times N}$, its Moore-Penrose pseudoinverse can be directly computed with an SVD-based method. Specifically, we first perform truncated SVD on $\mathcal{R}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\widetilde{\mathcal{T}}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I}\approx\mathbf{U}_{q}\Sigma_{q}\mathbf{V}_{q}^{\top}$, where $\mathbf{U}_{q}\in\mathbb{R}^{N\times q}$ and $\mathbf{V}_{q}\in\mathbb{R}^{N\times q}$ contains the first $q$ columns of $\mathbf{U}$ and $\mathbf{V}$. $\Sigma_{q}\in\mathbb{R}^{q\times q}$ is the diagonal matrix of $q$ largest singular values. It is a $q$-rank approximation of $\mathcal{R}$, which holds that $\mathrm{rank}(\mathrm{R})=q$. Then the Moore-Penrose pseudoinverse of $\mathcal{R}$ can be easily computed as follows:
$\mathcal{R}^{\dagger}=\mathbf{U}_{q}\Sigma_{q}^{-1}\mathbf{V}_{q}^{\top}.$ (17)
To leverage sparsity of $\mathcal{R}$ to avoid $\mathcal{O}(N^{3})$ complexity, we adopt the randomized SVD algorithm proposed by (Halko et al., 2011; Cai et al., 2023) to first approximate the range of the input matrix with a low-rank orthonormal matrix, and then perform SVD on this smaller matrix:
$\hat{\bm{U}}_{q},\hat{\Sigma}_{q},\hat{\bm{V}}_{q}^{\top}=\operatorname{ApproxSVD}(\mathcal{R},q),\quad\hat{\mathcal{R}}_{SVD}=\hat{\bm{U}}_{q}\hat{\Sigma}_{q}\hat{\bm{V}}_{q}^{\top},$ (18)
where $\hat{\bm{U}}_{q}$, $\hat{\Sigma}_{q}$, and $\hat{\bm{V}}_{q}$ are the approximated versions of $\mathbf{U}_{q}$, $\Sigma_{q}$, and $\mathbf{V}_{q}$. Then the Moore-Penrose pseudoinverse of $\mathcal{R}$ can be computed by:
$\mathcal{R}^{\dagger}=\hat{\bm{U}}_{q}\hat{\Sigma}^{-1}_{q}\hat{\bm{V}}_{q}^{\top}.$ (19)
The computation cost of randomized truncated SVD takes $\mathcal{O}(qK)$, where $K$ is the number of non-zero elements in $\mathcal{R}$, so we have $K=|E|$. Thus, the sparsity degree of $\mathcal{R}$ can determine the time complexity of its Moore-Penrose pseudoinverse, which demonstrates the importance of Lemma 4.1.
Appendix B Pseudo Code for CGNN
@@
The time complexity of randomized truncated SVD to compute $\mathcal{R}^{\dagger}$ is $\mathcal{O}(q|E|)$, and the message passing iteration has the same time complexity as GraphSAGE with $\mathcal{O}(|E|)$. Therefore, the overall time complexity of CGNN is $\mathcal{O}(q|E|)$.
Appendix C Implementation Details
C.1 Experimental Settings
We evaluate the performance by node classification accuracy with standard deviation in the semi-supervised setting. For Squirrel and Chameleon, we use 10 public splits (48%/32%/20% for training/validation/testing) provided by (Pei et al., 2019). For the remaining datasets, we adopt the same splits as (Tong et al., 2020a, 2021), which chooses 20 nodes per class for the training set, 500 for the validation set, and allocates the rest to the test set. We conduct our experiments on 2 Intel Xeon Gold 5215 CPUs and 1 NVIDIA GeForce RTX 3090 GPU.
C.2 Data Statistics
The dataset used in Section 5 are Squirrel, Chameleon (Rozemberczki et al., 2021), Citeseer (Sen et al., 2008), CoraML (Bojchevski & Günnemann, 2017), and AM-Photo (Shchur et al., 2018). We summarize their statistics in Table 4. $\mathrm{homo\_ratio}$ represents the homophily ratio, a metric proposed by Zhu et al. (2020). which is employed to gauge the degree of homophily within the graph. A lower $\mathrm{homo\_ratio}$ signifies a greater degree of heterophily, indicating a higher prevalence of edges that connect nodes of differing classes.

### v2

@@ 5.2 Similarity-based Graph Rewiring
Both the fundamental matrix defined in Eq. 3 and Weighted DiLap requires 4.1 to ensure the existence and uniqueness of the Perron vector $\pi$, conditions that are not universally met in general graphs. To fulfill the irreducibility and aperiodicity assumptions,  [3] introduce a teleporting probability uniformly distributed across all nodes. This method, inspired by PageRank [27], amends the transition matrix to $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, where $\gamma\in(0,1)$. $\mathbf{P}_{pr}$ allows for the possibility that a random walker may choose a non-neighbor node for the next step with a probability of $\frac{1-\gamma}{N}$. This adjustment ensures that $\mathbf{P}_{pr}$ is irreducible and aperiodic, so it has a unique $\pi$. However, this approach leads to a complete graph represented by a dense matrix $\mathbf{P}_{pr}$, posing significant challenges for subsequent computational processes.
@@
This rewiring approach satisfies 4.1 and maintains graph sparsity. Additionally, adding edges between nodes with similar features only minimally alters the overall semantics of the original graph. Based on $\widetilde{G}$ and its corresponding $\widetilde{\mathbf{P}}$ and $\widetilde{\mathbf{\Pi}}$, we have the deterministic Weighted DiLap $\widetilde{\mathcal{T}}$.
@@ 5.3 From DiLap to Commute Time
The proof is given in Section 8.2. Leveraging Lemma 5.1 and using Eq. 4, we can further compute the hitting times and commute times in terms of $\widetilde{\mathcal{T}}$ with the following theorem.
@@
Then we can derive the matrix forms of the hitting time $\mathcal{H}$ and commute time $\mathcal{C}$ as per Eq. 9:
@@
where $\odot$ denotes Hadamard product, and $\otimes$ is outer product. $\mathcal{H}_{ij}$ and $\mathcal{C}_{ij}$ correspond to the hitting and commute time from $v_{i}$ to $v_{j}$ respectively. The computation of commute times via DiLap, in contrast to the method delineated in Theorem 4.2, is primarily motivated by efficiency concerns. Specifically, Eq. 3 necessitates the inversion of a dense matrix with complexity $\mathcal{O}(N^{3})$, whereas our DiLap-based method hinges on computing the pseudoinverse of a sparse matrix $\mathcal{R}$. The pseudoinverse of $\mathcal{R}$ can be efficiently determined using SVD. Given the sparse nature of $\mathcal{R}$, we can employ well-established techniques such as the randomized truncated SVD algorithm [28, 29], which takes advantage of sparsity, to reduce the time complexity to $\mathcal{O}(q|E|)$, where $|E|$ denotes the number of edges reflecting the sparsity (See Section 8.3). Next, we present Commute Graph Neural Networks (CGNN) based on $\mathcal{C}$.
@@ 5.4 CGNN
where $\mathrm{Agg}_{\text{in}}^{(\ell)}(\cdot)$ and $\mathrm{Agg}_{\text{out}}^{(\ell)}(\cdot)$ are mean aggregation functions with different feature transformation weights, and $\mathrm{Comb}^{(\ell)}(\cdot)$ is a mean operator. Within each layer, the influence of $v_{j}$ on the central node $v_{i}$ is modulated by the commute-time-based proximity $\widetilde{\mathcal{C}}$ based on the edge directionality. We present the pseudocode of CGNN in Algorithm 1.
@@
The randomized truncated SVD to compute $\mathcal{R}^{\dagger}$ is $\mathcal{O}(q|E|)$ where $q$ is the required rank, and the message passing iteration has the same time complexity as DirGNN with $\mathcal{O}(L|E|d^{2})$. Therefore, the overall time complexity of CGNN is $\mathcal{O}((Ld^{2}+q)|E|)$. In practice, $q$ is typically set to 5, rendering the time complexity effectively linear with respect to the number of edges $|E|$. In GNN domain, models with a complexity less than $\mathcal{O}(N^{2})$ are generally considered feasible by researchers [30]. Given that real-world networks are often extremely sparse, i.e., $|E|\ll N^{2}$, CGNN demonstrates its feasibility as a model within the GNN family.
6 Experiments
6.1 Experimental Settings
We conduct node classification experiments on eight digraph datasets. We provide a performance comparison with 12 baselines including 1) General GNNs: GCN [31], GAT [32], and GraphSAGE [17]; 2) Non-local GNNs: APPNP [33], MixHop [34], GPRGNN [35], and GCNII [36]; 3) Digraph NNs: DGCN [4], DiGCN [3], MagNet [6], DiGCL [5], DUPLEX [19], and DirGNN [7]. For all baselines, we apply both the symmetrized and asymmetric adjacency matrix for experiments. The results reported are the better of the two results.
We evaluate the performance by node classification accuracy with standard deviation in the semi-supervised setting. For Squirrel and Chameleon, we use 10 public splits (48%/32%/20% for training/validation/testing) provided by [37]. For the remaining datasets, we adopt the same splits as [3, 5], which chooses 20 nodes per class for the training set, 500 for the validation set, and allocates the rest to the test set. We conduct our experiments on 2 Intel Xeon Gold 5215 CPUs and 1 NVIDIA GeForce RTX 3090 GPU.
6.2 Data Statistics
The datasets used in Section 6 are Squirrel, Chameleon [38], Citeseer [39], CoraML [40], AM-Photo [41], Snap-Patents, Roman-Empire, and Arxiv-Year [7]. We summarize their statistics in Table 1. $\mathrm{homo\_ratio}$ represents the homophily ratio, a metric proposed by [42]. which is employed to gauge the degree of homophily within the graph. A lower $\mathrm{homo\_ratio}$ signifies a greater degree of heterophily, indicating a higher prevalence of edges that connect nodes of differing classes.